# Enhancing Lip Reading with Multi-Scale Video and Multi-Encoder

### v1

@@ Abstract
Automatic lip-reading (ALR) aims to automatically transcribe spoken content from a speaker’s silent lip motion captured in video. Current mainstream lip-reading approaches only use a single visual encoder to model input videos of a single scale. In this paper, we propose to enhance lip-reading by incorporating multi-scale video data and multi-encoder. Specifically, we first propose a novel multi-scale lip extraction algorithm based on the size of the speaker’s face and an enhanced ResNet3D visual front-end (VFE) to extract lip features at different scales. For the multi-encoder, in addition to the mainstream Transformer and Conformer, we also incorporate the recently proposed Branchformer and E-Branchformer as visual encoders. In the experiments, we explore the influence of different video data scales and encoders on ALR system performance and fuse the texts transcribed by all ALR systems using recognizer output voting error reduction (ROVER). Finally, our proposed approach placed second in the ICME 2024 ChatCLR Challenge Task 2, with a 21.52% reduction in character error rate (CER) compared to the official baseline on the evaluation set.
@@
Fig. 1: Examples of multi-scale lip motion data of S217 (top) and S443 (bottom) speaker from the ChatCLR training set.
@@
Many researches have been conducted on ALR or audio-visual speech recognition (AVSR), which integrates visual features into the ASR system. In the early stage, Transformer [2], with its powerful sequence modeling capabilities, is widely used as the backbone for visual modeling in many studies [3, 4, 5]. However, as the Conformer [6] proposed, it has gradually replaced Transformer and become the mainstream choice for visual backbone, no matter in ALR [7, 8] or AVSR [9, 10] studies. Despite the surge in research on ALR or AVSR in recent years, most studies have focused on English, with less research dedicated to Chinese lip-reading. The absence of a large dataset for Chinese lip-reading was undoubtedly a significant reason, leaving researchers without a benchmark for validation and fair comparison. Since 2021, the Multimodal Information Based Speech Processing (MISP) Challenge series [11, 12, 13] has been successfully held for three consecutive editions to advance the research on Chinese audio-visual speech processing and released a large and challenging Chinese audio-visual dataset [11, 14]. In addition, the hosting of the Chinese Continuous Visual Speech Recognition Challenge (CNVSRC) 2023  [^1] http://cnceleb.org/competition further promotes research in Chinese lip-reading, releasing a larger volume (over 300 hours) of high-quality Chinese audio-visual data [15]. However, it mainly focuses on professional announcers, speechmakers, or bloggers with prepared topics, limiting some practical applicability.
@@
To promote universal Chinese lip-reading in real-world scenarios, the ICME 2024 Chat-scenario Chinese Lipreading (ChatCLR) Challenge  [^2] https://mispchallenge.github.io/ICME2024 is launched. The target speaker lip-reading track (Task 2) is designed for lip-reading in the free talk scenarios, and over 100 hours of far-field real-recorded video data are released. This paper elaborates on our proposed lip-reading approach that builds and fuses multiple ALR systems with different visual encoders and multi-scale lip motion video data. For data extraction, we design 6 different lip regions of interest (ROIs) according to the size of the speaker’s face and the center of the lip. Based on the works of Wang et al. [16, 17], we propose an Enhanced ResNet3D visual front-end (VFE) module for better visual feature extraction. Moreover, recently proposed Branchformer [18] and E-Branchformer [19] encoders are introduced as the backbone for visual modeling, constructing more diverse ALR systems with different scales of video data. Finally, the recognizer output voting error reduction [20] (ROVER) technique is used to fuse the transcripts from all ALR systems. Overall, our contributions are as follows:
@@
  • We propose an algorithm to extract multi-scale lip motion videos based on the size of the speaker’s face.
@@
  • We proposed an Enhanced ResNet3D visual front-end module based on the ResNet [21] and 3D convolutions.
@@
  • We systematically compare different visual encoders and scales of video data on the performance of ALR.
@@ 2.1 Multi-Scale Lip Video Data Extraction
Many researches [22, 17] have indicated the performance of lip-reading recognition systems is not solely determined by the motion of the speaker’s lips. Video data from the cheeks, chin, and even forehead can improve recognition accuracy. Moreover, the winning approach proposed by Wang et al. [17] in CNVSRC 2023 also shows the performance of ALR systems improves with an increase in the field of view of the video data centered around the lips. However, their method of enlarging the lip ROI by a fixed adjustment of the crop size for each video frame (from 48 to 112) does not consider the discrepancy among all speakers’ face sizes or the distances between the camera and speakers. This could lead to inconsistencies in the facial area covered by the obtained lip motion videos, even when using the same crop size.
Therefore, we design a lip motion extraction method based on the size of the speaker’s face in the video. The official ChatCLR Challenge provides the face and lip coordinates on the image in almost every frame of all video data. For a video sequence of speaker S with a total number of frames $T$ (25 frames per second), we define $t$ as the number of frames where the speaker’s face is successfully detected ($t\leq T$). The coordinates of the top-left and bottom-right corners of the $i$-th detected face are defined as $(Lface^{i}_{x},Lface^{i}_{y})$ and $(Rface^{i}_{x},Rface^{i}_{y})$, respectively. Thus, the crop size $L$, that is, the side of the square lip ROI, can be formulated as:
@@
In cases where the speaker’s face is obscured or not directed towards the camera, it becomes hard to detect the position of their lips. To overcome this, for a failed lip detection frame, we search for the nearest detected lip forward or backward in time. This allows us to calculate the center coordinates for every frame and extract lip motion video data as accurately as possible. It is worth noting that if the detection rate of the face or lip in a segment of the video does not exceed 50%, it will be discarded. All lip motion data extracted at any scale will be resized to a uniform size of $112\times 112$.
Figure 1 shows the lip motion data extracted from two speakers’ videos under different scaling factors. It can be observed that when the $scale$ is set to 0.6 or 0.8, the extracted lip motion data almost only contains the speaker’s lips. As the value of $scale$ increases, areas such as the chin, cheeks, nose, and eyes gradually enter the lip-centered ROI.
@@ 3.1 Data Processing
Dataset. All ALR systems in this paper are built with the training and development dataset released by the ChatCLR Challenge Task 2. The training, development, and evaluation sets include 110.95, 4.5, and 2.41 hours, respectively, all free talk video data, each recorded by 2 to 6 speakers in the TV living room. The recording equipment is a 1080P resolution camera with a 120-degree wide-angle lens, placed 3 to 5 meters away from the speakers, allowing it to capture every speaker simultaneously. For all recorded video data, the challenge official provides the face and lip coordinates of almost every speaker in each frame. In addition, all datasets provide ground-truth timestamps for the speakers. The training and development sets also offer text labels.
@@ 3.2 Implementation Details
Model Configuration. All systems are built and decoded using the open-source toolkit Espnet [24]. To build diverse ALR systems, we use different scales of lip motion videos and visual encoders during training and inference. For the Branchformer encoder, we set the hidden size to 256, the number of encoder layers to 24, the dropout rate to 0.2, the cgmlp linear units to 2048, and the convolutional kernels to 31. The E-Branchformer includes 12 encoder layers, with the rest of the settings being the same as those for the Branchformer. To compare fairly with the current mainstream Transformer and Conformer encoders, we adjust their parameters to ensure that different encoders have similar parameter sizes. Specifically, we use a 24-layer Transformer and an 11-layer Conformer, both with a 2048 feed-forward dimension. Other settings are consistent with the Branchformer. All systems use the Enhanced ResNet3D visual front-end for video feature extraction. Four ResNet3D blocks have 3, 4, 6, and 3 basic blocks, respectively, with the feature dimensions of 32, 54, 128, and 256. 6-layer Transformers, with 256 hidden sizes and a 0.2 dropout rate, serve as decoders in all ALR systems.
Training Strategy. Each ALR system is trained from scratch for 50 epochs using the Adam optimizer with a learning rate of 1e-3, weight decay of 1e-6, and a warm-up strategy for the first 5 epochs. The learning rate increased linearly with the number of training steps. After the training, we select the top 20 epochs with the lowest development loss for averaging as the decoding model on the development set and the initial model for the next fine-tuning stage. We use the same speakers’ data in the training set as in the evaluation set, 90% of the development set data as the fine-tuning training data, and the remaining 10% as the fine-tuning development set. Set the learning rate to 5e-4 with the Adam optimizer, linearly warm-up for 5 epochs, and fine-tune for 20 epochs. Finally, we average the top 10 epochs with the lowest loss on the development set as the decoding model for the evaluation set.
@@
System Encoder Scale Param Dev $\downarrow$ Eval $\downarrow$
S1 Branchformer 0.6 52.6M 81.97 85.87
S2 Branchformer 0.8 52.6M 80.35 84.69
S3 Transformer 1.0 53.3M 80.95 85.56
S4 Conformer 50.8M 81.82 84.97
S5 E-Branchformer 50.0M 79.02 83.44
S6 Branchformer 52.6M 78.86 83.40
S7 Branchformer 1.25 52.6M 77.27 82.12
S8 Transformer 1.5 53.3M 79.13 84.97
S9 Conformer 50.8M 79.88 83.51
S10 E-Branchformer 50.0M 77.06 82.45
S11 Branchformer 52.6M 76.21 81.56
S12 E-Branchformer 1.75 50.0M 76.01 82.61
S13 Branchformer 52.6M 75.80 81.95
R1 ROVER (S3 - S6) 75.32 80.71
R2 ROVER (S8 - S11) 73.49 79.58
R3 ROVER (S1, S2, S6, S7, S11, S13) 73.45 79.58
R4 ROVER ALL 72.02 78.17
@@ 3.3.1 Which visual encoder performs best?
We conduct comparative experiments at two video data scales, 1.0 and 1.5, to compare the influence of different visual encoders on the performance of the ALR system. Table 1 shows detailed experimental results. At the 1.0 scale, the ALR systems based on Branchformer (S6) and E-Branchformer (S5) encoders perform similarly, with Branchformer showing only a 0.04% improvement in CER on the evaluation set. However, compared to mainstream Transformer (S3) and Conformer (S4) encoders, Branchformer achieves noteworthy improvements of 2.16% and 1.57%, respectively. At the 1.5 scale, Branchformer (S11) ’s advantage is more pronounced, with reductions of 1.05%, 1.95%, and 3.41% in CER on the evaluation set compared to E-Branchformer (S10), Conformer (S9), and Transformer (S8), respectively. Combining the experiments at both data scales, the ALR system based on the Branchformer encoder shows the best performance, followed by the E-Branchformer, showing a big improvement over the Transformer and Conformer.
@@ 3.3.2 Which data scale is most suitable?
As analyzed in the previous section, the ALR system based on the Branchformer encoder performs the best. Therefore, we use the Branchformer as the visual encoder for our data scale comparison experiments, conducted at all 6 data scales: 0.6, 0.8, 1.0, 1.25, 1.5, and 1.75. The results are shown in Table 1. It can be observed that within the range of 0.6 to 1.5 data scales, the larger the data scale used for model training and inference, the better the performance of the ALR system. The system (S11) based on the Branchformer encoder built with 1.5-scale data shows a 4.31% reduction in CER on the evaluation set compared to the one (S1) using 0.6-scale data. However, as the data scale increases from 1.5 to 1.75, the ALR systems based on Branchformer (S13) and E-Branchformer (S12) encoders see their evaluation set CERs increase by 0.39% and 0.16%, respectively. In summary, the ALR system built with lip motion data at a scale of 1.5, under the same visual encoder, performs the best.
3.3.3 How much yield does multi-system fusion bring?
To roundly show the benefits of multi-system fusion based on ROVER, we report not only the results of fusing all ALR systems (R4) but also the fusion of ALR systems built with different encoders (R1, R2) and those using the same encoder but different scale data (R3), as shown in Table 1. First, we fuse ALR systems built with 1.0-scale video data based on four visual encoders. The fused result R1 shows a 2.69% reduction in CER on the evaluation set compared to the best-performing ALR system (S6) at the 1.0 scale. Similarly, the result fusing all ALR systems built with 1.5 scale data (R2) reduces CER by 1.98% compared to the best-performing system (S11) at the 1.5 scale. In addition, the result of fusing all ALR systems based on the Branchformer encoder but built with different scale data (R3) also shows a 1.98% CER reduction compared to the best one (S11). These results sufficiently demonstrate that significant improvements can be achieved through text fusion using ROVER, whether for ALR systems built with different visual encoders or data scales. Finally, the R4 system that fused all ALR systems achieves a CER of 78.17% on the evaluation set, which is a 3.39% reduction compared to the best-performing system S11, ranking second place in the ICME 2024 ChatCLR Challenge Task2.
@@ 3.3.3 How much yield does multi-system fusion bring?
System Method Param Dev $\downarrow$ Eval $\downarrow$
A0 ResNet3D Baseline 47.3M 83.19 88.93
A1 Enhanced ResNet3D 52.6M 81.23 87.45
A2 + Data Augmentation 52.6M 76.27 83.84
A3    + Speed Perturb 52.6M 76.21 82.44
S11       + Fine-tune 52.6M 76.21 81.56

### v2

@@ Abstract
Automatic lip-reading (ALR) aims to automatically transcribe spoken content from a speaker’s silent lip motion captured in video. Current mainstream lip-reading approaches only use a single visual encoder to model input videos of a single scale. In this paper, we propose to enhance lip-reading by incorporating multi-scale video data and multi-encoder. Specifically, we first propose a novel multi-scale lip motion extraction algorithm based on the size of the speaker’s face and an Enhanced ResNet3D visual front-end (VFE) to extract lip features at different scales. For the multi-encoder, in addition to the mainstream Transformer and Conformer, we also incorporate the recently proposed Branchformer and E-Branchformer as visual encoders. In the experiments, we explore the influence of different video data scales and encoders on ALR system performance and fuse the texts transcribed by all ALR systems using recognizer output voting error reduction (ROVER). Finally, our proposed approach placed second in the ICME 2024 ChatCLR Challenge Task 2, with a 21.52% reduction in character error rate (CER) compared to the official baseline on the evaluation set.
@@
Fig. 1: Examples of multi-scale lip motion videos of speaker S217 (top) and S443 (bottom) from the ChatCLR training set.
@@
Many researches have been conducted on ALR or audio-visual speech recognition (AVSR), which integrates visual features into the ASR system. In the early stage, Transformer [2], with its powerful sequence modeling capabilities, is widely used as the backbone for visual modeling in many studies [3, 4, 5]. However, as the Conformer [6] proposed, it has gradually replaced Transformer and become the mainstream choice for visual backbone, no matter in ALR [7, 8] or AVSR [9, 10] studies. Despite the surge in research on ALR or AVSR in recent years, most studies have focused on English, with less research dedicated to Chinese lip-reading. The absence of a large dataset for Chinese lip-reading was undoubtedly a significant reason, leaving researchers without a benchmark for validation and fair comparison. Since 2021, the Multimodal Information based Speech Processing (MISP) Challenge series [11, 12, 13] has been successfully held for three consecutive editions to advance the research on Chinese audio-visual speech processing and released a large and challenging Chinese audio-visual dataset [11, 14]. In addition, the hosting of the Chinese Continuous Visual Speech Recognition Challenge (CNVSRC) 2023  [^1] http://cnceleb.org/competition further promotes research in Chinese lip-reading, releasing a larger volume (over 300 hours) of Chinese audio-visual dataset [15]. However, it mainly focuses on professional announcers, speechmakers, or bloggers with prepared topics, limiting some practical applicability.
@@
To promote universal Chinese lip-reading in real-world scenarios, the ICME 2024 Chat-scenario Chinese Lipreading (ChatCLR) Challenge  [^2] https://mispchallenge.github.io/ICME2024 is launched. The target speaker lip-reading track (Task 2) is designed for lip-reading in the free talk scenarios, and over 100 hours of far-field real-recorded video data are released. This paper elaborates on our proposed lip-reading approach that builds and fuses multiple ALR systems with different visual encoders and multi-scale lip motion video data. For data extraction, we design 6 different lip regions of interest (ROIs) according to the size of the speaker’s face and the coordinate of the lip. Based on our previous works [16, 17], we propose an Enhanced ResNet3D visual front-end (VFE) module for better visual feature extraction. Moreover, recently proposed Branchformer [18] and E-Branchformer [19] encoders are introduced as the backbone for visual modeling, constructing more diverse ALR systems with different scales of video data. Finally, the recognizer output voting error reduction [20] (ROVER) technique is used to fuse the transcripts from all ALR systems. Overall, our contributions are as follows:
@@
  • We introduce an algorithm to extract multi-scale lip motion videos based on the size of the speaker’s face.
@@
  • We propose an Enhanced ResNet3D visual front-end module based on the ResNet [21] and 3D convolutions.
@@
  • We systematically compare different visual encoders and video data scales on the performance of ALR.
@@ 2.1 Multi-Scale Lip Video Data Extraction
Many researches [22, 17] have indicated the performance of lip-reading recognition systems is not solely determined by the motion of the speaker’s lips. Visual information from the cheeks, chin, and even forehead can improve recognition accuracy. Moreover, the winning approach proposed by Wang et al. [17] in CNVSRC 2023 also shows the performance of ALR systems improves with an increase in the field of view of the video data centered around the lips. However, their method of enlarging the lip ROI by a fixed adjustment of the crop size for each video frame (from 48 to 112) does not consider the discrepancy among all speakers’ face sizes or the distances between the camera and speakers. This could lead to inconsistencies in the facial area covered by the obtained lip motion videos, even when using the same crop size.
Therefore, we design a lip motion extraction method based on the size of the speaker’s face in the video. The official ChatCLR dataset provides all speakers’ face and lip coordinates on the image in almost every video frame. For a video sequence of speaker S with a total number of frames $T$ (25 frames per second), we define $t$ as the number of frames where the speaker’s face and lip are successfully detected ($t\leq T$). The coordinates of the top-left and bottom-right corners of the $i$-th detected face are defined as $(Lface^{i}_{x},Lface^{i}_{y})$ and $(Rface^{i}_{x},Rface^{i}_{y})$, respectively. Thus, the crop size $L$, that is, the side of the square lip ROI, can be formulated as:
@@
In cases where the speaker’s face is obscured or not directed towards the camera, it becomes hard to detect their lips. To overcome this, for a failed lip detection frame, we search for the nearest detected lip forward or backward in time. This allows us to calculate the center coordinates for every frame and extract lip motion video data as accurately as possible. It is worth noting that if the detection rate of the face or lip in a segment of the video does not exceed 50%, it will be discarded. All lip motion data extracted at any scale will be resized to a uniform size of $112\times 112$.
Figure 1 shows the lip motion video extracted from two speakers under different scaling factors. It can be observed that when the $scale$ is set to 0.6 or 0.8, the extracted lip motion data almost only contains the speaker’s lips. As the value of $scale$ increases, areas such as the chin, cheeks, nose, and eyes gradually enter the lip-centered ROI.
@@ 3.1 Data Processing
Dataset. All ALR systems in this paper are built with the training and development dataset released by the ChatCLR Challenge Task 2. The training, development, and evaluation sets include 110.95, 4.5, and 2.41 hours, respectively, all free talk video data, each recorded by 2 to 6 speakers in the TV living room. The recording equipment is a 1080P resolution camera with a 120-degree wide-angle lens, placed 3 to 5 meters away from the speakers, allowing it to capture every speaker simultaneously. For all recorded video data, the challenge official provides speakers’ face and lip coordinates in almost every frame.
@@ 3.2 Implementation Details
Model Configuration. All ALR systems are built based on the open-source toolkit Espnet [24]. To build diverse ALR systems, we use different scales of lip motion videos and visual encoders during training and inference. For the Branchformer encoder, we set the hidden size to 256, the number of encoder layers to 24, the dropout rate to 0.2, the cgmlp linear units to 2048, and the convolutional kernels to 31. The E-Branchformer includes 12 encoder layers, with the rest of the settings being the same as those for the Branchformer. To compare fairly with the current mainstream Transformer and Conformer encoders, we adjust their parameters to ensure that different encoders have similar parameter sizes. Specifically, we use a 24-layer Transformer and an 11-layer Conformer, both with a 2048 feed-forward dimension. All systems use the Enhanced ResNet3D visual front-end for feature extraction. Four ResNet3D blocks have 3, 4, 6, and 3 basic blocks, respectively, with the feature dimensions of 32, 54, 128, and 256. Six-layer Transformers, with 256 hidden sizes and a 0.2 dropout rate, serve as decoders in all ALR systems.
Training Strategy. Each ALR system is trained for 50 epochs using the Adam optimizer with a learning rate of 1e-3, weight decay of 1e-6, and a linear warm-up strategy for the first 5 epochs. After the training, we average the top 20 epoch models with the lowest development loss as the decoding model on the development set and the initial model for the fine-tuning stage. For fine-tuning, we use the same speakers’ data in the training set as in the evaluation set and 90% of the development set data as the training data, while the remaining 10% as the fine-tuning development set. Set the learning rate to 5e-4 with the Adam optimizer, linearly warm-up for 5 epochs, and fine-tune for 20 epochs. Finally, we average the top 10 epochs with the lowest loss on the fine-tuning development set as the decoding model for the evaluation set.
@@
System Encoder Scale Param (M) Dev $\downarrow$ Eval $\downarrow$
S1 Transformer 0.6 53.3 83.09 87.61
S2 Conformer 50.8 83.61 87.14
S3 E-Branchformer 50.0 82.21 86.29
S4 Branchformer 52.6 81.97 85.87
S5 Branchformer 0.8 52.6 80.35 84.69
S6 Transformer 1.0 53.3 80.95 85.56
S7 Conformer 50.8 81.82 84.97
S8 E-Branchformer 50.0 79.02 83.44
S9 Branchformer 52.6 78.86 83.40
S10 Branchformer 1.25 52.6 77.27 82.12
S11 Transformer 1.5 53.3 79.13 84.97
S12 Conformer 50.8 79.88 83.51
S13 E-Branchformer 50.0 77.06 82.45
S14 Branchformer 52.6 76.21 81.56
S15 E-Branchformer 1.75 50.0 76.01 82.61
S16 Branchformer 52.6 75.80 81.95
R1 ROVER (S1 - S4) 78.78 83.15
R2 ROVER (S6 - S9) 75.32 80.71
R3 ROVER (S11 - S14) 73.49 79.58
R4 ROVER (S4, S5, S9, S10, S14, S16) 73.45 79.58
R5 ROVER ALL 72.02 78.17
@@ 3.3.1 Which visual encoder performs best?
To explore the influence of different visual encoders on the performance of the ALR system, we conduct comparative experiments at 3 video data scales, 0.6, 1.0, and 1.5. Table 1 shows detailed experimental results. At the 0.6 and 1.0 scales, the ALR systems based on Branchformer (S4, S9) and E-Branchformer (S3, S8) encoders perform similarly, with Branchformers showing only 0.42% and 0.04% improvements in CER on the evaluation set at 0.6 and 1.0 scale, respectively. However, compared to mainstream Transformers (S1, S6) and Conformers (S2, S7), Branchformer achieves noteworthy improvements of 1.74% and 1.27% at 0.6 scale, while 2.16% and 1.57% at 1.0 scale. At the 1.5 scale, Branchformer (S14) ’s advantage is more pronounced, with reductions of 1.05%, 1.95%, and 3.41% in CER on the evaluation set compared to E-Branchformer (S13), Conformer (S12), and Transformer (S11), respectively. In summary, the ALR system based on the Branchformer encoder shows the best performance, followed by the E-Branchformer, showing a big improvement over the Transformer and Conformer.
@@ 3.3.2 Which data scale is most suitable?
As analyzed in the previous section, the ALR system based on the Branchformer encoder performs the best. Therefore, we use the Branchformer as the visual encoder for our data scale comparison experiments, conducted at all 6 data scales: 0.6, 0.8, 1.0, 1.25, 1.5, and 1.75. The results are shown in Table 1. It can be observed that within the range of 0.6 to 1.5 data scales, the larger the data scale used for model training and inference, the better the performance of the ALR system. The system (S14) based on the Branchformer encoder built with 1.5-scale data shows a 4.31% reduction in CER on the evaluation set compared to the one (S4) using 0.6-scale data. However, as the data scale increases from 1.5 to 1.75, the ALR systems based on Branchformer (S16) and E-Branchformer (S15) encoders see their evaluation set CERs increase by 0.39% and 0.16%, respectively. In summary, the ALR system built with lip motion video data at a scale of 1.5, under the same visual encoder, performs the best.
3.3.3 How much gain does multi-system fusion bring?
To roundly show the benefits of multi-system fusion based on ROVER, we report not only the results of fusing all ALR systems (R5) but also the fusion of ALR systems built with different encoders (R1, R2, R3) and those using the same encoder but different scale data (R4), as shown in Table 1. First, we fuse ALR systems built with 0.6-scale video data based on four visual encoders. The fused result R1 shows a 2.72% reduction in CER on the evaluation set compared to the best-performing ALR system (S4) at the 0.6 scale. Similarly, the results fusing all ALR systems built with 1.0 (R2) and 1.5 scale data (R3) reduce CER by 2.69% and 1.98%, compared to the best-performing systems at the 1.0 scale (S9) and 1.5 scale (S14), respectively. In addition, the result of fusing all ALR systems based on the Branchformer encoder but built with different scale data (R4) also shows a 1.98% CER reduction compared to the best one (S14). These results sufficiently demonstrate that significant improvements can be achieved through text fusion using ROVER, whether for ALR systems built with different visual encoders or data scales. Finally, the result R5 that fused all ALR systems achieves a CER of 78.17% on the evaluation set, which is a 3.39% reduction compared to the best-performing system (S14), ranking second place in the ICME 2024 ChatCLR Challenge Task2.
@@ 3.3.3 How much gain does multi-system fusion bring?
System Method Param (M) Dev $\downarrow$ Eval $\downarrow$
A0 ResNet3D Baseline 47.3 83.19 88.93
A1 Enhanced ResNet3D 52.6 81.23 87.45
A2 + Data Augmentation 52.6 76.27 83.84
A3      + Speed Perturb 52.6 76.21 82.44
S14         + Fine-tune 52.6 76.21 81.56